### v1

@@ Abstract
Chemical cartography of the Galactic disk provides insights to its structure and assembly history over cosmic time. In this work, we use chemical cartography to explore chemical gradients and azimuthal substructure in the Milky Way disk with giant stars from APOGEE DR17. We confirm the existence of a radial metallicity gradient in the disk of [Fe/H]/R dex/kpc and a vertical metallicity gradient of [Fe/H]/Z . We find azimuthal variations ( dex) on top of the radial metallicity gradient that have been previously established with other surveys. The APOGEE giants show strong correlations with stellar age and the intensity of azimuthal variations in [Fe/H]; older stellar populations show the largest deviations from the radial metallicity gradient. Beyond iron, we show that other elements (e.g., Mg, O) display azimuthal variations at the dex-level across the Galactic disk. We illustrate that moving into the orbit-space could help constrain the mechanisms producing these azimuthal metallicity variations. These results suggest that the spiral arms of the Galaxy are not solely responsible for azimuthal metallicity variations and other Galactic processes are at play.
@@ 1 Introduction
There are two distinct mechanisms that spiral arms can generate to induce azimuthal metallicity variations. The first process is radial migration/churning (Sellwood & Binney, 2002) which, in this context, is the morphing of stellar orbits that alters the angular momentum of an orbit without the addition of excess energy. The second process is kinematic heating/blurring which would induce changes (heat) in the stellar orbital parameters without increasing the angular momentum. An example of this is the spiral arms dynamically evoking changes in the motions of stars along the leading and trailing edges of the spiral arms (Grand et al., 2012, 2016).
@@
Using Gaia DR3 stars, Poggio et al. (2022) found that the azimuthal variations present in the disk correlate with where the spiral arms are predicted to be using different samples of bright stars (using effective temperature as a proxy for age) within kpc of the Sun. Interestingly, they concluded the inverse of the simulated results from Di Matteo et al. (2013), in that the youngest population exhibits the most pronounced inhomogenities.
@@ 2 Data
The initial sample of stars came from the seventeenth data release (DR17, Abdurro’uf et al., 2022) of the Apache Point Observatory Galactic Evolution Experiment (APOGEE) (Majewski et al., 2017b). APOGEE is a large scale near-infrared (15140Å < < 16960Å) stellar spectroscopic survey. The survey spans both hemispheres, consisting of a spectrograph on the 2.5-meter Sloan Foundation Telescope at the Apache Point Observatory in New Mexico, USA as well as the 2.5-meter Irénée du Pont Telescope at the Las Campanas Observatory in Chile.
We made use of the astroNN value-added catalog (Leung & Bovy, 2019a)111The astroNN package is available here: https://github.com/henrysky/astroNN of abundances, distances, and ages for APOGEE DR17 sources. astroNN is an open source python package developed for the neural network trained on the APOGEE data and is designed to be a general package for deep learning in astronomy. The APOGEE-astroNN catalog contains results from applying astroNN neural nets on APOGEE DR17 spectra to infer stellar parameters, abundances trained with ASPCAP DR17 (Holtzman et al., in prep.), distances retrained with Gaia eDR3 (Gaia Collaboration et al., 2021) from Leung & Bovy (2019b), and ages trained with APOKASC-2 (Mackereth et al., 2019) in combination with low-metallicity asteroseismic ages (Montalbán et al., 2021).
@@ 2.1 Planar Thin Disk Sample
The density of stars in the stellar disk of the Milky Way is thought to noticeably decline at about kpc away from the Galactic center (Momany et al., 2006; Reylé et al., 2009; Carraro et al., 2010). The inner radii of the Galactic disk are thought to be dominated by the bar, typically believed to extend out to kpc (Hammersley et al., 1994; Wegg et al., 2015; Lucey et al., 2023). Thus to minimize the chemical and dynamical effects driven by the bar in the most central region, and the drop in the density of stars in the furthest regions, we adopted a Galactocentric radius cut of kpc to our final thin disk sample. We place a 0.3 kpc cut on Z to ensure that we are only selecting bona fide thin disk stars with orbits confined to the Galactic plane. After these final selection cuts, we have a planar thin disk sample of 32,768 stars with the median uncertainties for Teff, log , and [Fe/H] being 33K, 0.07 dex, and 0.03 dex, respectively.
@@ 4.2 Azimuthal Variations in [Fe/H]/R
There is a blue-red-blue-red pattern in [Fe/H] shown in the final panel of Figure 5. The red bins represent areas of the Milky Way that are more metal-rich than the model predicts and the blue bins are where the model over-estimates the stellar metallicities. This oscillating pattern has been observed previously with different stellar surveys, such as Gaia DR3 stars in Figure 2 of Poggio et al. (2022), with similar results to this work. Hawkins (2022) use LAMOST OBAF stars as well as Gaia DR3 stars, uncovering azimuthal variations in both of these populations, showing similar patterns in which the solar-neighborhood is more metal-rich (red) and the inner/outer galaxy are more metal-poor (blue).
@@
The black solid lines in the bottom panel of Figure 6 represent the spiral arms as determined by Reid et al. (2019) using the parallaxes and proper motions associated with high-mass star-forming regions using a Very Long Baseline Array, European VLBI Newtork, and the Japanese VLBI Exploration of Radio Astrometry project. They locate multiple arm segments with pitch angles ranging from 7∘ to 20∘ with the widths of the arms increasing with distance from the Galactic center. The black contours in the top panel of Figure 6 represent the spiral arm locations determined by Poggio et al. (2021) using the over-density of upper main sequence stars determined by Gaia DR3. These results were consistent with some of the Reid et al. (2019) arms, such as the Sagittarius-Carina spiral arm, while the geometry of arms with Galactic longitudes from 180∘ to 270∘ were significantly different from other spiral arm models. The metal-rich and metal-poor portions of the Milky Way in our work are not fully encompassed by either determination of the spiral arm locations, thus we look towards other processes that may have induced this azimuthal substructure.
If the spiral arm locations do not obviously coincide with the azimuthal substructure, this begs the question, what is causing these variations? Carr et al. (2022) quantified the effect that a Sagittarius (Sgr)-like dwarf galaxy would have on the kinematics and chemsitry of the Milky Way-like disk upon first interaction. They did this by painting the particles in their N-body simulation of the disk with an negative radial [Fe/H] gradient. The passage of Sgr through the plane will cause radial rearrangement and disrupt stellar orbits. If the heating of orbits is due to a non-axisymmetric feature, such as Sgr, migration and mixing signatures will manifest as an approximate quadrupole in chemical azimuthal variations across the disk. In their present day snapshot of a simulated Milky Way being perturbed by a Sgr-like dwarf galaxy, Carr et al. (2022) found that the maximum departure from the [Fe/H]/R gradient occurs in the solar annulus on the side of disk closest to the current position of Sgr, while the minimum is found adjacent to Sgr in the outermost annuli. These azimuthal variations in metallicity will be dependent on the model of Sgr chosen. It is interesting to note that in our sample the solar annulus appears to be more metal-rich, although we lack a complete view of the Galaxy to classify this definitively.
@@ 4.4 Azimuthal Metallicity Variations by Age
If the azimuthal variations are exclusively natal in origin (i.e., they are a result of variations in the gas-phase abundances at the time a population was born), then we would predict that younger populations would exhibit strong metallicity variations due to these populations forming more recently. However, if azimuthal variations are exclusively the result of dynamical processes, we would predict that older populations would show stronger azimuthal variations because these stars would have had more time to interact with non-axisymmetric features in the Galaxy that would cause the rearrangement of stars, leading to the observed azimuthal metallicity variations. To distinguish between these predictions, we applied the present-day ages from APOGEE-astroNN to separate our sample by stellar age.
We separated the APOGEE giants into 3 distinct age groups to explore the effects that age has on departures from the metallicity gradient. The age bins we selected are 0-2 Gyr, 2-6 Gyr, and Gyr, with 8730, 18493 and 5545 stars, respectively. The strongest radial metallicity gradient appears in the middle age group with [Fe/H]/R dex/kpc, outranking the youngest ([Fe/H]/R dex/kpc) and oldest ([Fe/H]/R dex/kpc) age groups.
@@
After following the same methodology as Section 4.2 and subtracting the data from the linear gradient model, Figure 8 shows the [Fe/H] residuals for each of the 3 age groups. Visually, the signatures seem to be the strongest in the panel containing the oldest stars, while the deviations lessen in intensity with decreasing age.
In Figure 9 we plot azimtuhal angle on the X axis and metallicity excess on the Y axis for each of the different age bins. We show that the oldest age group has the most extreme variations in terms of [Fe/H], with the youngest stars showing minimal azimuthal variations. Interestingly, there seems to be a phase offset in the curves among the different age bins. This phase offset is also seen in Figure 12 of Carr et al. (2022). For all age groups, the variations are minimized at , which is most likely due to the sample selection and the number of stars located along the line of sight from the Sun towards the center of the Galaxy. The future releases of SDSS V will provide better radial and azimuthal coverage to alleviate this problem.
We find that the strongest azimuthal [Fe/H] variations are apparent in solar-age stars. Additionally, the metallicity excess is largest in the oldest stars (6+ Gyr) and similar in the youngest (0-2 Gyr) and intermediate aged (2-6 Gyr) stars, with a phase offset that differentiates between them. These results are in line with the simulated results of Debattista et al. (2024). The authors found that azimuthal metallicity variations were still present when excluding younger populations, whereas we find the strongest azimuthal metallicity variations in our older populations. Although azimuthal vartiations have been quantified in younger populations (e.g., the OBA sample in Hawkins, 2022), the presence of the strongest azimuthal variations in our older populations suggest an important contribution from dynamical processes to the creation of azimuthal metallicity variations throughout the Galactic disk. Thus, in the following section, we delve into possible dynamical origins.
@@ 4.5 Linking Chemistry to Dynamics
Figure 10 is a six panel plot showing the aforementioned parameters versus . The bins are colored by density with the pink stars representing the running medians in each panels with error bars showing the standard deviation. There is a slight positive trend with and Jr, suggesting that stars with a larger metallicity excess tend to have higher radial actions. This correlation with radial action follows the predictions of the Debattista et al. (2024) simulation222Note that those authors suggest that Jr should be computed as a time-averaged quantity. We do not investigate the impact of time-averaging, and leave further investigation of this to future work.. There are no obvious trends with the angular momentum, vertical action, maximum height, and total energy of stars when comparing to .
@@
In the first and third panels of Figure 10, we see larger at higher eccentricities and radial actions. This implies that blurring (heating of orbits) contributes a non-negligible amount to the mechanisms that are causing these observed azimuthal metallicity variations. In reality, multiple Galactic mechanisms are likely responsible for the azimuthal metallicity variations and more work is required to pin down mechanism responsible for the variations observed in this work and others (e.g., Hawkins, 2022; Poggio et al., 2022; Imig et al., 2023).
@@ 5 Summary
Azimuthal variations are found throughout the disk in [Fe/H]. These deviations were quantified by modelling the linear gradient and subtracting the model from the data to illustrate the [Fe/H]-poor to [Fe/H]-rich oscillating azimuthal pattern in the disk. These results are consistent with other azimuthal studies that have used different datasets or tracer populations (Eilers et al., 2022; Hawkins, 2022). While previous studies have found a correlation with azimuthal metallicity variations and the location of the spiral arms, we find some deviations to that in this work. We argue this is worth further investigation over a wider range of azimuthal angles (Hackshaw et al. in prep) (Section 4.2).
@@
Azimuthal substructure is seen in elements beyond iron in [/Fe]. While [Fe/H] exhibits the strongest variations from the gradient, the -elements (O, Mg, Si, S) still show azimuthal variations on the dex-level. These -abundances manifest as anti-correlations with [Fe/H], following the predictions by the production mechanisms between and Fe-peak elements (Section 4.3).
@@
Azimuthal substructure varies by stellar age. Older populations (6+ Gyr) exhibit more extreme deviations from the radial metallicity gradient than seen in younger (0-2 Gyr) and intermediate (2-6) populations, which suggests that the azimuthal metallicity variations largely stem from dynamical processes rather than natal processes. This age dependence can be coupled with mock observed simulations to identify why the older populations are disproportionately affected by the cause of these azimuthal variations (Section 4.4).

### v2

@@ Abstract
Chemical cartography of the Galactic disk provides insights to its structure and assembly history over cosmic time. In this work, we use chemical cartography to explore chemical gradients and azimuthal substructure in the Milky Way disk with giant stars from APOGEE DR17. We confirm the existence of a radial metallicity gradient in the disk of [Fe/H]/R dex/kpc and a vertical metallicity gradient of [Fe/H]/Z . We find azimuthal variations ( dex) on top of the radial metallicity gradient that have been previously established with other surveys. The APOGEE giants show strong correlations with stellar age and the intensity of azimuthal variations in [Fe/H]; young populations and intermediate-aged populations both show significant deviations from the radial metallicity gradient. Beyond iron, we show that other elements (e.g., Mg, O) display azimuthal variations at the dex-level across the Galactic disk. We illustrate that moving into the orbit-space could help constrain the mechanisms producing these azimuthal metallicity variations in the future. These results suggest that dynamical processes play an important role in the formation of azimuthal metallicity variations.
@@ 1 Introduction
Azimuthal metallicity variations can be generated by different mechanisms in different stellar populations. In older populations, dynamical processes are predicted to be the most important. In younger stars, dynamical processes are thought to have less of an effect on their orbits due to these stars interacting less with features in the Galaxy that can alter their orbits. If azimuthal metallicity variations were present in these younger populations, they would trace any variations in the gas from which they were born (Spitoni et al., 2019).
There are two distinct dynamical mechanisms that spiral arms can generate to induce azimuthal metallicity variations. The first process is radial migration/churning (Sellwood & Binney, 2002) which, in this context, is the morphing of stellar orbits that alters the angular momentum of an orbit without the addition of excess energy. The second process is kinematic heating/blurring which would induce changes (heat) in the stellar orbital parameters without increasing the angular momentum. An example of this is the spiral arms dynamically evoking changes in the motions of stars along the leading and trailing edges of the spiral arms (Grand et al., 2012, 2016).
@@
Using Gaia DR3 stars, Poggio et al. (2022) found that the azimuthal variations present in the disk correlate with where the spiral arms are predicted to be using different samples of bright stars (using effective temperature as a proxy for age) within kpc of the Sun. When considering older stellar populations, Di Matteo et al. (2013) provides context illustrating that the strength of the azimuthal variations is highly dependent on the strength of the perturbations.
@@ 2 Data
The initial sample of stars came from the seventeenth data release (DR17, Abdurro’uf et al., 2022) of the Apache Point Observatory Galactic Evolution Experiment (APOGEE) (Majewski et al., 2017b). APOGEE is a large scale near-infrared (15140Å < < 16960Å) stellar spectroscopic survey. The survey spans both hemispheres, consisting of a spectrograph (Wilson et al., 2019) on the 2.5-meter Sloan Foundation Telescope at the Apache Point Observatory in New Mexico (Gunn et al., 2006), USA as well as the 2.5-meter Irénée du Pont Telescope at the Las Campanas Observatory in Chile (Bowen & Vaughan, 1973). Detailed explanation of the data processing and reduction can be found in Nidever et al. (2015).
We made use of the astroNN value-added catalog (Leung & Bovy, 2019a)111The astroNN package is available here: https://github.com/henrysky/astroNN of abundances, distances, and ages for APOGEE DR17 sources. astroNN is an open source python package developed for the neural network trained on the APOGEE data and is designed to be a general package for deep learning in astronomy. The APOGEE-astroNN catalog contains results from applying astroNN neural nets on APOGEE DR17 spectra to infer stellar parameters, abundances trained with ASPCAP DR17 (Holtzman et al., in prep.; García Pérez et al., 2016; Shetrone et al., 2015; Smith et al., 2021), distances retrained with Gaia eDR3 (Gaia Collaboration et al., 2021) from Leung & Bovy (2019b), and ages trained with APOKASC-2 (Mackereth et al., 2019) in combination with low-metallicity asteroseismic ages (Montalbán et al., 2021).
@@ 2.1 Planar Thin Disk Sample
The density of stars in the stellar disk of the Milky Way is thought to noticeably decline at about kpc away from the Galactic center (Momany et al., 2006; Reylé et al., 2009; Carraro et al., 2010). The inner radii of the Galactic disk are thought to be dominated by the bar, typically believed to extend out to kpc (Hammersley et al., 1994; Wegg et al., 2015; Lucey et al., 2023). Thus to minimize the chemical and dynamical effects driven by the bar in the most central region, and the drop in the density of stars in the furthest regions, we adopted a Galactocentric radius cut of kpc to our final thin disk sample. We place a 0.3 kpc cut on Z to ensure that we are only selecting bona fide thin disk stars with orbits confined to the Galactic plane. Finally, we remove any stars in clusters as identified by Hunt & Reffert (2023) which reduced our sample by stars. After these final selection cuts, we have a planar thin disk sample of 25,404 stars with the median uncertainties for Teff, log , and [Fe/H] being 33K, 0.07 dex, and 0.03 dex, respectively.
@@ 4.2 Azimuthal Variations in [Fe/H]/R
There is a blue-red-blue-red pattern in [Fe/H] shown in the final panel of Figure 5. The red bins represent areas of the Milky Way that are more metal-rich than the model predicts and the blue bins are where the model over-estimates the stellar metallicities. These signatures remain the same when different cuts of Zmax were tested. This oscillating pattern has been observed previously with different stellar surveys, such as Gaia DR3 stars in Figure 2 of Poggio et al. (2022), with similar results to this work. Hawkins (2022) use LAMOST OBAF stars as well as Gaia DR3 stars, uncovering azimuthal variations in both of these populations, showing similar patterns in which the solar-neighborhood is more metal-rich (red) and the inner/outer galaxy are more metal-poor (blue). We explored whether the APOGEE selection function has an impact on recovered azimuthal variations using the selection function from the gaiaunlimited package (Castro-Ginard et al., 2023). We find that azimuthal variations are recoverable even with the impact of APOGEE’s selection function.
@@
The black solid lines in the bottom panel of Figure 6 represent the spiral arms as determined by Reid et al. (2019) using the parallaxes and proper motions associated with high-mass star-forming regions using a Very Long Baseline Array, European VLBI Network, and the Japanese VLBI Exploration of Radio Astrometry project. They locate multiple arm segments with pitch angles ranging from 7∘ to 20∘ with the widths of the arms increasing with distance from the Galactic center. The black contours in the top panel of Figure 6 represent the spiral arm locations determined by Poggio et al. (2021) using the over-density of upper main sequence stars determined by Gaia DR3. These results were consistent with some of the Reid et al. (2019) arms, such as the Sagittarius-Carina spiral arm, while the geometry of arms with Galactic longitudes from 180∘ to 270∘ were significantly different from other spiral arm models. The metal-rich and metal-poor portions of the Milky Way in our work are not fully encompassed by either determination of the spiral arm locations, thus we look towards other processes that may have induced this azimuthal substructure.
Carr et al. (2022) quantified the effect that a Sagittarius (Sgr)-like dwarf galaxy would have on the kinematics and chemistry of the Milky Way-like disk upon first interaction. They did this by painting the particles in their N-body simulation of the disk with an negative radial [Fe/H] gradient. The passage of Sgr through the plane will cause radial rearrangement and disrupt stellar orbits. If the heating of orbits is due to a non-axisymmetric feature, such as Sgr, migration and mixing signatures will manifest as an approximate quadrupole in chemical azimuthal variations across the disk. In their present day snapshot of a simulated Milky Way being perturbed by a Sgr-like dwarf galaxy, Carr et al. (2022) found that the maximum departure from the [Fe/H]/R gradient occurs in the solar annulus on the side of disk closest to the current position of Sgr, while the minimum is found adjacent to Sgr in the outermost annuli. These azimuthal variations in metallicity will be dependent on the model of Sgr chosen. It is interesting to note that in our sample the solar annulus appears to be more metal-rich, although we lack a complete view of the Galaxy to classify this definitively.
@@ 4.4 Azimuthal Metallicity Variations by Age
If the azimuthal variations are exclusively natal in origin (i.e., they are a result of variations in the gas-phase abundances at the time a population was born), then we would predict that younger populations would exhibit strong metallicity variations due to these populations forming more recently. To examine the effect that stellar age has on the strength of azimuthal metallicity variations, we applied the present-day ages from APOGEE-astroNN to separate our sample by stellar age.
We separated the APOGEE giants into 3 distinct age groups to explore the effects that age has on departures from the metallicity gradient. The age bins we selected are 0-2 Gyr, 2-6 Gyr, and Gyr, with 7297, 14419 and 3890 stars, respectively. The strongest radial metallicity gradient appears in the middle age group with [Fe/H]/R dex/kpc, outranking the youngest ([Fe/H]/R dex/kpc) and oldest ([Fe/H]/R dex/kpc) age groups. We note that the radial metallicity trend is not well represented by a linear gradient in the oldest age bin.
@@
After following the same methodology as Section 4.2 and subtracting the data from the linear gradient model, Figure 8 shows the [Fe/H] residuals for each of the 3 age groups. Visually, the signatures seem to be the strongest in the panel containing the oldest stars, while the deviations lessen in intensity with decreasing age. Our methodology relies on the assumption that the radial metallicity gradient is linear in each age bin. Since the oldest population is not well-represented by a linear radial metallicity gradient, we do not draw any assumptions about the azimuthal structure for stars with ages of Gyr. This population is included in Figures 8 and 9 for completeness.
In Figure 9 we plot azimtuhal angle on the X axis and metallicity excess on the Y axis for each of the different age bins. We show that the most azimuthal variations are present in both young and intermediate age stars. There may be substructure in the oldest age bin, but these ages are not well-sampled azimuthally thus we refrain from drawing robust conclusions. For all age groups, the variations are minimized at sin, which is most likely due to the sample selection and the number of stars located along the line of sight from the Sun towards the center of the Galaxy. The future releases of SDSS V will provide better radial and azimuthal coverage to alleviate this problem.
We find that the strongest azimuthal [Fe/H] variations are apparent in solar-age stars. These results are in line with the simulated results of Debattista et al. (2024). The authors found that azimuthal metallicity variations were still present when excluding younger populations, whereas we find the strongest azimuthal metallicity variations in our intermediate and older populations. Although azimuthal variations have been quantified in younger populations (e.g., the OBA sample in Hawkins, 2022), the presence of the strongest azimuthal variations in our older populations suggest an important contribution from dynamical processes to the creation of azimuthal metallicity variations throughout the Galactic disk. Thus, in the following section, we delve into possible dynamical origins.
@@ 4.5 Linking Chemistry to Dynamics
Figure 10 is a six panel plot showing the aforementioned parameters versus . The bins are colored by density with the pink stars representing the running medians in each panels with error bars showing the standard error in each eccentricity bin. There is a slight positive trend with and Jr, suggesting that stars with a larger metallicity excess tend to have higher radial actions. This correlation with radial action follows the predictions of the Debattista et al. (2024) simulation222Note that those authors suggest that Jr should be computed as a time-averaged quantity. We do not investigate the impact of time-averaging, and leave further investigation of this to future work.. There are no obvious trends with the angular momentum, vertical action, maximum height, and total energy of stars when comparing to . To test if the older stars are solely responsible for the trends we are seeing in the dynamical properties, we redid this analysis for both our young and intermediate aged populations. These trends still hold in both age groups, thus we conclude that these dynamical trends are not just a byproduct of the oldest stars.
@@
In the first and third panels of Figure 10, we see larger at higher eccentricities and radial actions. This could imply that blurring (heating of orbits) contributes a non-negligible amount to the mechanisms that are causing these observed azimuthal metallicity variations. In reality, multiple Galactic mechanisms are likely responsible for the azimuthal metallicity variations and more work is required to pin down mechanism responsible for the variations observed in this work and others (e.g., Hawkins, 2022; Poggio et al., 2022; Imig et al., 2023).
@@ 5 Summary
Azimuthal variations are found throughout the disk in [Fe/H]. These deviations were quantified by modelling the linear gradient and subtracting the model from the data to illustrate the [Fe/H]-poor to [Fe/H]-rich oscillating azimuthal pattern in the disk. These results are consistent with other azimuthal studies that have used different datasets or tracer populations (Poggio et al., 2022; Hawkins, 2022). While previous studies have found a correlation with azimuthal metallicity variations and the location of the spiral arms, we find some deviations to that in this work. We argue this is worth further investigation over a wider range of azimuthal angles (Hackshaw et al. in prep) (Section 4.2).
@@
Azimuthal substructure is seen in elements beyond iron in [/Fe]. While [Fe/H] exhibits the strongest variations from the gradient, the -elements (O, Mg, Si, S) still show azimuthal variations on the dex-level. These -abundances manifest as anti-correlations with [Fe/H] in some sections of the Galactic disk, following the predictions by the production mechanisms between and Fe-peak elements (Section 4.3).
@@
Azimuthal substructure varies by stellar age. Azimuthal variations are quantified in both the young (0-2 Gyr) and intermediate (2-6 Gyr) populations. These variations in the young population suggests that the origins of this substructure could be either natal or dynamical, while the azimuthal variations that are present in the intermediate age bin suggests that these variations likely stem from dynamical processes rather than natal processes. The statistics are too low in the oldest age bin to draw robust conclusions. This age dependence can be coupled with mock observed simulations to identify the possible dynamical origins of these azimuthal variations (Section 4.4).